OpenStudy (aonz):

y=2x Is this Odd, Even or neither? Pls explain

OpenStudy (anonymous):

A function is even if the following is true for the function: \[f(-x)=f(x).\] Visually, you'd see symmetry across the y-axis if you were to graph the function. A function is odd if \[f(-x)=-f(x).\] Here, the function would be symmetric about the origin.

OpenStudy (aonz):



OpenStudy (aonz):

wouldnt this be odd?

OpenStudy (aonz):

f(−x)= -2x −f(x)= -2x

OpenStudy (anonymous):

Yes

OpenStudy (aonz):

in the back of my book it says neither...

OpenStudy (aonz):

maybe the back of book is wrong?

OpenStudy (anonymous):

It's got to be; wolfram seems to agree http://www.wolframalpha.com/input/?i=is+y%3D2x+odd%3F

OpenStudy (anonymous):

And it's quite clear to me that it is indeed an odd function.

OpenStudy (aonz):

k thanks